3. Consider the following table for Canadas GDP in 2016 Components of GDP in 2016 (billions of dollars) Category Personal expenditure on consumer goods and services Durable goods 231.2 273.4 648.9 28.6 Non-durable goods Services Non-profit institutions serving households final consumption expenditure General governments final consumption expenditure Business gross fixed capital formation 429.8 Residential structures 154.1 Non-residential structures Machinery and equipment 80.2 33.3 2.8 Intellectual property products Non-profit institutions serving households gross fixed capital formation General governments gross fixed capital formation 78.9 2.2 628.7 676.7 Investment in inventories Exports of goods and services Imports of goods and services a. Calculate 2016 consumer spending b. Calculate 2016 private investment spending c. Calculate 2016 net exports d. Calculate 2016 government purchases of goods and services and government investment spending e. Calculate 2016 gross domestic produdt f. Calculate 2016 consumer spending on services as a percentage of total consumer spending g. Calculate 2016 exports as a percentage of imports. h. Calculate 2016 exports and imports as a percentage of gross domestic product. Was there a trade surplus or trade deficit in 2016? How large was this as a percentage of GDP?
Answer #1

Consumer Spending 2016

Durable Goods 231.2

Non Durable Goods 273.4

Services 648.9

Non profit Institutions Final Consumption Expenditure = 28.6

Adding all Above

Consumer Spending 2016 = 1182.1 Billion Dollars

2. Private Investment Spending

Non Residential Structures 117.1

Residential Structures 154.1

Machinery and Equipment 80.2

Intellectual Property Products 33.3

Investment in Inventories -2.2

Non profit Institutions serving Household Gross Fixed Capital Formation 2.8

Adding All Above

Private Investment Expenditure = 385.3 Billion Dollars

3. Net Exports

Export of Goods and Services - Imports of Goods and Services

628.7 - 676.7 = - 48 Billion Dollars

4.Government Purchase of Goods and Services and Government Investment Spending

It is equal to:

General Government Final Consumption Expenditure + General Government Gross Fixed Capital Formation

= 429.8 + 78.9 = 508.7 Billion Dollars